# InputDSA : Demixing then comparing recurrent and externally driven dynamics

### v1

@@ 2.3 Estimating Linear Operators
Here $x_{t}$ represents the measured state of the system at time $t$, $\phi$ is a nonlinear embedding of the data that typically expands the dimensionality of the state space, and AA is a matrix that is identified using some variant of least-squares regression. The goal of the Dynamic Mode Decomposition is to approximate the Koopman Operator (Koopman, 1931), a theoretical object that exists for all dynamical systems which encodes the linear dynamics of observables (functions that act on the state) under the system dynamics. Prior work has explored many different choices of $\phi$. For example, $\phi$ can be a kernel function, a delay embedding, or even a neural network (Williams et al., 2016; Brunton et al., 2017; Arbabi & Mezić, 2017; Lusch et al., 2018). Intuitively, the dimensionality expansion acts similarly to the kernel trick (Smola & Schölkopf, 1998), where embedding into higher dimensions ‘unfolds’ the nonlinearity. The DMD can be applied in non-autonomous systems, although this risks mixing driving and intrinsic dynamics (Proctor et al., 2016a).
@@
While estimating AA and $B$ via DMDc is an intuitive extension to input-driven systems, it has a hidden failure mode in the analysis of partially-observed systems. This is particularly important in the analysis of neural data, in which a small subset of neurons in a vast population are recorded. Generically, an input-driven system that is partially observed receives input to both the observed and unobserved components. The input at time $t$ therefore affects the observed state at time $t$ (instantaneously) and in future time steps through the unobserved state (Fig. 2A). This means that simply applying DMDc in this setting will result in the $B$ matrix becoming biased toward the intrinsic dynamics of the system. We develop a formal description of this problem in Appendix D. We solve this problem by introducing Subspace DMDc, a novel extension of Subspace DMD (Takeishi et al., 2017b) that incorporates input. In brief, Subspace DMDc utilizes subspace identification algorithms from classical control theory (Verhaegen & Verdult, 2007), which seek to identify linear dynamical systems of the form:
@@
With only $y_{t}$ and $u_{t}$ observed. The situation of partial observability is a special case of this problem. In practice, we use the well-known N4SID or PO-MOESP algorithms to estimate AA and $B$ (Van Overschee & De Moor, 1994; Verhaegen, 1994) on lifted states (thereby leveraging the power of nonlinear DMD algorithms such as Williams et al. 2016, although). For further technical details on the subspace identification algorithm, see Appendix Section E. We also highlight that subspace identification methods are designed to handle data with both observation and process noise, therefore providing InputDSA noise robustness (Verhaegen, 1994; Verhaegen & Verdult, 2007).
@@ 4.2 InputDSA captures differences in neural population dynamics across time
Figure 5: InputDSA quantifies differences in neural population dynamics across task epoch. (A) Auditory evidence accumulation task schematic (adapted from Luo et al. 2025). (B) Trial-averaged neural trajectories visualized in the top two Principal Components. Stars indicate a ”neural time of commitment” (nTc): the time point when the curvature of trial-averaged trajectories is maximum (marked by stars). (C) Similarity of neural dynamics before and after the nTc for rat T223. Bars denote standard error across 21 sessions. (D) Distribution of top real eigenvalues of state-transition matrix AA and fit power law for pre- vs. post-commitment activity. Left, sample distribution. Right, distribution of power law exponents across sessions. Dots denote individual sessions, lines indicate paired periods within session, likewise in E and F. (E) Normalized effects of intrinsic and input-driven dynamics in pre vs. post periods. (F) Subspace angles of the input and state operators within and between time periods. $A_{pre}-A_{pre}$ (likewise $B_{pre}$) denotes the noise floor via split-halves comparison.
Comparing neural dynamics before and after the nTc (“Pre” vs. “Post”), we found significant shifts in both intrinsic and input-driven dynamics, consistent with the changes at nTc reported in (Luo et al., 2025) (Fig. 5C). To probe how the intrinsic dynamics change, we analyzed the eigenspectrum of the state-transition matrix AA estimated by SubspaceDMDc before and after the nTc. Each spectrum was fit with a power law $\lambda_{i}\;\propto\;\left(\frac{i}{N}\right)^{\alpha},$ where $\lambda_{i}$ are the $i$-th eigenvalue sorted in descending order. We found that the post-commitment periods consistently showed smaller $\alpha$, indicating slower decay and thus longer-lasting intrinsic dynamics (Fig. 5D). This is directly related to the controllability of the dynamical system, which describes how easy it is for an input sequence to drive the system to arbitrary points in state space (Luenberger, 1979). Smaller DMD eigenvalues implies greater input controllability, which would be expected for a more input-driven system as Luo et al. (2025) identified is the case in the pre-nTc regime.
Likewise, the average magnitude of the intrinsic dynamics strengthen while the input-driven dynamics weaken in the Post-nTC period, reflecting a transition into more autonomous, less input-sensitive regime after the nTc (Fig. 5E). By measuring the subspace angle between the AA and $B$ operators across the nTc, we find that the neural dynamics occupy different subspaces across the nTc, reflected a reorganization of dynamics at the decision-commitment time (Fig. 5F). Together, these results suggest that population activity undergoes a regime shift at the nTc: transitioning from an input-driven, evidence-accumulation phase into an intrinsically dominated, decision-commitment phase, as suggested by (Luo et al., 2025).
@@ Appendix B Dynamic Mode Decomposition with Control (DMDc)
Dynamic Mode Decomposition with control (DMDc) (Proctor et al., 2016a) extends standard DMD to dynamical systems with external inputs. It provides a data-driven approximation of both the intrinsic dynamics AA and input couplings $B$, enabling system identification and forecasting for non-autonomous dynamical systems. Here, we briefly review the formulation of DMDc. For full details, please refer to Proctor et al. (2016a). In practice, we can apply DMDc whenever the state is fully observed. When this is not the case, refer to Sections D and E.
@@
We can further project AA and $B$ into the system’s state space using
@@ B.1 On nonlinear embeddings in DMDc
In the standard DMDc formulation (above), an SVD is taken across $\Omega$, which concatenates the state data $X$ with the input data $U$. Although this has the benefit of whitening across all regressors, it can bias the estimation of AA and $B$ depending on the relative scalings and dimensionalities of $X$ and $U$. This has a critical effect when applying high-dimensional nonlinear embeddings to only $X$ ($U$) individually, as the SVD will be increasingly dominated by signal from $X$ ($U$) if the data is sufficiently rich. Therefore, whenever we apply delay embeddings or other nonlinear embeddings to $X$, we do so commensurately to $U$.
@@ Appendix D Partial Observation Induces Biases in Input Operator B
These equations show that when performing regression as in DMDcDMDc on partially-observed, delay-embedded data, the estimates of $B$ become biased by the intrinsic dynamics in the unobserved states. Biases in $B$ emerges when utilizing delay embeddings as dimensionality expansions, as we can see from the above formulation. Although we display the formal connection with linear systems above, it is simple to observe that the same problem occurs with nonlinear dynamics as well.
@@ E.2 N4SID
Numerical Algorithms for Subspace State Space System Identification (N4SID) (Van Overschee & De Moor, 1994) utilizes a similar approach as the above to jointly estimate $A,B,C,D$ operators in a state space model from data $Y$ and $U$. Here we briefly describe the algorithm that we apply to estimate AA and $B$ that are used for comparison of partially observed systems, as first defined by Overschee & Moor (1994). We used code from https://github.com/spmvg/nfoursid/tree/master for our implementation of n4sid. For the Subspace DMDc, we lift to a nonlinear space before state estimation.
@@ Appendix G Solving for optimal orthogonal $C$ efficiently
Where the metric is once again Eq. 38. This improves the optimization stability on AA until at least dimension $150$ for O(0.001) error per element error, which is more than enough in practice for InputDSA . We have the following lemma which states that this metric captures equivalency between two linear systems.
@@
In practice, taking large matrices AA to many powers results in numerical instability issues, especially when $\lambda_{max}(A)>1$. Algorithmically, we check the condition number of $A^{n}$ before choosing to include the term in the controllability matrix. If it is too small or too large, we stop.
@@ Appendix I Hyperparameter tuning for InputDSA
SubspaceDMDc involves one rank parameter $r$, corresponding to the dimensionality of the latent state space. In practice, selecting an $r$ slightly higher than the true state dimension often yields a better estimation of the AA matrix.
@@
    A value $\text{MASE}<1$ indicates that DMDc predicts next-step activity (using the estimated operators AA and $B$) more accurately than simply copying the current time step.
  • Model complexity: The estimated operators AA and $B$ should not be overly complex or dominated by spurious features (e.g., many small eigenvalues clustered near zero). To assess this, we compute the Akaike Information Criterion (AIC) for next-step prediction on the test set. AIC balances predictive accuracy against model complexity and, in our setting, is given by
@@
where AA is the drift amplitude (noise level) and $f$ is a low drift rate.
@@
Appendix L Ordinary Least Squares Biases Estimates of AA in the presence of input noise
Using two datasets: the nonlinear dynamical systems as in Fig. 2, and the RNNs trained on Random Target Task as in Fig. 3, we show that increasing the noise variance systematically contracts the singular spectrum of $B$ toward zero. Because the true input effect is underfit, the SubspaceDMDc regression inflates the real part of AA’s eigenvalues to absorb the variance in the inputs that correlates with the state.
Figure 11: The eigenspectrum of the AA operator and singular spectrum of the $B$ operator when the input time-series is corrupted by Gaussian or pink noise of different variance.

### v2

@@ 2.3 Estimating Linear Operators
Here $x_{t}$ represents the measured state of the system at time $t$, $\phi$ is a nonlinear embedding of the data that typically expands the dimensionality of the state space, and $A$ is a matrix that is identified using some variant of least-squares regression. The goal of the Dynamic Mode Decomposition is to approximate the Koopman Operator (Koopman, 1931), a theoretical object that exists for all dynamical systems which encodes the linear dynamics of observables (functions that act on the state) under the system dynamics. Prior work has explored many different choices of $\phi$. For example, $\phi$ can be a kernel function, a delay embedding, or even a neural network (Williams et al., 2016; Brunton et al., 2017; Arbabi & Mezić, 2017; Lusch et al., 2018). Intuitively, the dimensionality expansion acts similarly to the kernel trick (Smola & Schölkopf, 1998), where embedding into higher dimensions ‘unfolds’ the nonlinearity. The DMD can be applied in non-autonomous systems, although this risks mixing driving and intrinsic dynamics (Proctor et al., 2016a).
@@
While estimating $A$ and $B$ via DMDc is an intuitive extension to input-driven systems, it has a hidden failure mode in the analysis of partially-observed systems. This is particularly important in the analysis of neural data, in which a small subset of neurons in a vast population are recorded. Generically, an input-driven system that is partially observed receives input to both the observed and unobserved components. The input at time $t$ therefore affects the observed state at time $t$ (instantaneously) and in future time steps through the unobserved state (Fig. 2A). This means that simply applying DMDc in this setting will result in the $B$ matrix becoming biased toward the intrinsic dynamics of the system. We develop a formal description of this problem in Appendix D. We solve this problem by introducing Subspace DMDc, a novel extension of Subspace DMD (Takeishi et al., 2017b) that incorporates input. In brief, Subspace DMDc utilizes subspace identification algorithms from classical control theory (Verhaegen & Verdult, 2007), which seek to identify linear dynamical systems of the form:
@@
With only $y_{t}$ and $u_{t}$ observed. The situation of partial observability is a special case of this problem. In practice, we use the well-known N4SID or PO-MOESP algorithms to estimate $A$ and $B$ (Van Overschee & De Moor, 1994; Verhaegen, 1994) on lifted states (thereby leveraging the power of nonlinear DMD algorithms such as Williams et al. 2016, although). For further technical details on the subspace identification algorithm, see Appendix Section E. We also highlight that subspace identification methods are designed to handle data with both observation and process noise, therefore providing InputDSA noise robustness (Verhaegen, 1994; Verhaegen & Verdult, 2007).
@@ 4.2 InputDSA captures differences in neural population dynamics across time
Figure 5: InputDSA quantifies differences in neural population dynamics across task epoch. (A) Auditory evidence accumulation task schematic (adapted from Luo et al. 2025). (B) Trial-averaged neural trajectories visualized in the top two Principal Components. Stars indicate a ”neural time of commitment” (nTc): the time point when the curvature of trial-averaged trajectories is maximum (marked by stars). (C) Similarity of neural dynamics before and after the nTc for rat T223. Bars denote standard error across 21 sessions. (D) Distribution of top real eigenvalues of state-transition matrix $A$ and fit power law for pre- vs. post-commitment activity. Left, sample distribution. Right, distribution of power law exponents across sessions. Dots denote individual sessions, lines indicate paired periods within session, likewise in E and F. (E) Normalized effects of intrinsic and input-driven dynamics in pre vs. post periods. (F) Subspace angles of the input and state operators within and between time periods. $A_{pre}-A_{pre}$ (likewise $B_{pre}$) denotes the noise floor via split-halves comparison.
Comparing neural dynamics before and after the nTc (“Pre” vs. “Post”), we found significant shifts in both intrinsic and input-driven dynamics, consistent with the changes at nTc reported in (Luo et al., 2025) (Fig. 5C). To probe how the intrinsic dynamics change, we analyzed the eigenspectrum of the state-transition matrix $A$ estimated by SubspaceDMDc before and after the nTc. Each spectrum was fit with a power law $\lambda_{i}\;\propto\;\left(\frac{i}{N}\right)^{\alpha},$ where $\lambda_{i}$ are the $i$-th eigenvalue sorted in descending order. We found that the post-commitment periods consistently showed smaller $\alpha$, indicating slower decay and thus longer-lasting intrinsic dynamics (Fig. 5D). This is directly related to the controllability of the dynamical system, which describes how easy it is for an input sequence to drive the system to arbitrary points in state space (Luenberger, 1979). Smaller DMD eigenvalues implies greater input controllability, which would be expected for a more input-driven system as Luo et al. (2025) identified is the case in the pre-nTc regime.
Likewise, the average magnitude of the intrinsic dynamics strengthen while the input-driven dynamics weaken in the Post-nTC period, reflecting a transition into more autonomous, less input-sensitive regime after the nTc (Fig. 5E). By measuring the subspace angle between the $A$ and $B$ operators across the nTc, we find that the neural dynamics occupy different subspaces across the nTc, reflected a reorganization of dynamics at the decision-commitment time (Fig. 5F). Together, these results suggest that population activity undergoes a regime shift at the nTc: transitioning from an input-driven, evidence-accumulation phase into an intrinsically dominated, decision-commitment phase, as suggested by (Luo et al., 2025).
@@ Appendix B Dynamic Mode Decomposition with Control (DMDc)
Dynamic Mode Decomposition with control (DMDc) (Proctor et al., 2016a) extends standard DMD to dynamical systems with external inputs. It provides a data-driven approximation of both the intrinsic dynamics $A$ and input couplings $B$, enabling system identification and forecasting for non-autonomous dynamical systems. Here, we briefly review the formulation of DMDc. For full details, please refer to Proctor et al. (2016a). In practice, we can apply DMDc whenever the state is fully observed. When this is not the case, refer to Sections D and E.
@@
We can further project $A$ and $B$ into the system’s state space using
@@ B.1 On nonlinear embeddings in DMDc
In the standard DMDc formulation (above), an SVD is taken across $\Omega$, which concatenates the state data $X$ with the input data $U$. Although this has the benefit of whitening across all regressors, it can bias the estimation of $A$ and $B$ depending on the relative scalings and dimensionalities of $X$ and $U$. This has a critical effect when applying high-dimensional nonlinear embeddings to only $X$ ($U$) individually, as the SVD will be increasingly dominated by signal from $X$ ($U$) if the data is sufficiently rich. Therefore, whenever we apply delay embeddings or other nonlinear embeddings to $X$, we do so commensurately to $U$.
@@ Appendix D Partial Observation Induces Biases in Input Operator B
These equations show that when performing regression as in $DMDc$ on partially-observed, delay-embedded data, the estimates of $B$ become biased by the intrinsic dynamics in the unobserved states. Biases in $B$ emerges when utilizing delay embeddings as dimensionality expansions, as we can see from the above formulation. Although we display the formal connection with linear systems above, it is simple to observe that the same problem occurs with nonlinear dynamics as well.
@@ E.2 N4SID
Numerical Algorithms for Subspace State Space System Identification (N4SID) (Van Overschee & De Moor, 1994) utilizes a similar approach as the above to jointly estimate $A,B,C,D$ operators in a state space model from data $Y$ and $U$. Here we briefly describe the algorithm that we apply to estimate $A$ and $B$ that are used for comparison of partially observed systems, as first defined by Overschee & Moor (1994). We used code from https://github.com/spmvg/nfoursid/tree/master for our implementation of n4sid. For the Subspace DMDc, we lift to a nonlinear space before state estimation.
@@ Appendix G Solving for optimal orthogonal $C$ efficiently
Where the metric is once again Eq. 38. This improves the optimization stability on $A$ until at least dimension $150$ for O(0.001) error per element error, which is more than enough in practice for InputDSA . We have the following lemma which states that this metric captures equivalency between two linear systems.
@@
In practice, taking large matrices $A$ to many powers results in numerical instability issues, especially when $\lambda_{max}(A)>1$. Algorithmically, we check the condition number of $A^{n}$ before choosing to include the term in the controllability matrix. If it is too small or too large, we stop.
@@ Appendix I Hyperparameter tuning for InputDSA
SubspaceDMDc involves one rank parameter $r$, corresponding to the dimensionality of the latent state space. In practice, selecting an $r$ slightly higher than the true state dimension often yields a better estimation of the $A$ matrix.
@@
    A value $\text{MASE}<1$ indicates that DMDc predicts next-step activity (using the estimated operators $A$ and $B$) more accurately than simply copying the current time step.
  • Model complexity: The estimated operators $A$ and $B$ should not be overly complex or dominated by spurious features (e.g., many small eigenvalues clustered near zero). To assess this, we compute the Akaike Information Criterion (AIC) for next-step prediction on the test set. AIC balances predictive accuracy against model complexity and, in our setting, is given by
@@
where $A$ is the drift amplitude (noise level) and $f$ is a low drift rate.
@@
Appendix L Ordinary Least Squares Biases Estimates of $A$ in the presence of input noise
Using two datasets: the nonlinear dynamical systems as in Fig. 2, and the RNNs trained on Random Target Task as in Fig. 3, we show that increasing the noise variance systematically contracts the singular spectrum of $B$ toward zero. Because the true input effect is underfit, the SubspaceDMDc regression inflates the real part of $A$’s eigenvalues to absorb the variance in the inputs that correlates with the state.
Figure 11: The eigenspectrum of the $A$ operator and singular spectrum of the $B$ operator when the input time-series is corrupted by Gaussian or pink noise of different variance.